FIND A PA모TICULAR 뛰0UTON ye FoR A31) At FINO 요 x222x
Answer #1

2) 3 eAt- T+ At At 3! ( Then 3 A - where 4 ( ( 3 3 A A = A 4 - 3 4 7 3 -/ A3- A A = 4 -3 - / ond erefont 3 5 3 e4t 4 -I -3 7+44= -- whch can be Jewriten as - + 4D+4) Dt tD +4) ex - 2A (Pta)2 (Dt2-2)2 eax D2 D (F-) lous tant f(Dra)